two numbers total 78 and have a difference of 6
Found 2 solutions by rfer, math-vortex:
Answer by rfer(16322)

x+x+6=78
2x=72
x=36
x+6=42

Answer by math-vortex(648)

Two numbers total 78 and have a difference of 6. I presume that you want to know what the two numbers are.
========================
.
Let
x = the first number
y = the second number
.
The sum of the two numbers is 78. Using algebra, we write the equation,
x + y = 78
.
The difference between the number is 6. Using algebra, we write,
x - y = 6
.
Now we have a system of two equations in two variables. The system lends itself nicely to solving by elimination. Add the two equations together.
.
x + y = 78
x - y = 6
----------
2x + 0 = 84
x = 42
.
Now we know that the first number is 42. Since the difference between the two numbers is 6, the second number is 42 - 6 = 36.
.
The sum of 42 and 36 is 78, so these are our two numbers.
.
--------------
I solved this using a system of linear equations. You can also solve it by writing an expression for the 2nd number in terms of x.
.
Let
x = 1st number
x - 6 = 2nd number
.
The sum of the two numbers is 78, so
x + (x-6) = 78
.
Now solve for x, and proceed as I did above.
.
hope this helps!
.
Ms.Figgy
math.in.the.vortex